OpenStudy (anonymous):

Find the quotient of the complex numbers. If necessary, use the slash bar ( / ) to enter a fraction. (1 - 7i )/(6 - 2i )

OpenStudy (anonymous):

help

OpenStudy (apoorvk):

Hmm, so we have something like \(\large \frac{z_1}{z_2}\), now multiply both the numerator the denominator by the conjugate of \(z_2\)

OpenStudy (anonymous):

the answer ?

OpenStudy (anonymous):

multiply numerator and denominator by the conjugate of the denoninator the conjugate of \(a+bi\) is \(a-bi\) and this works because \()a+bi)(a-bi)=a^2+b^2\) a real number

OpenStudy (anonymous):

is the answer .5-i ?

OpenStudy (anonymous):

denominator will be \(6^2+2^2=40\) numerator will be whatever you get when you multiply \[(1-7i)(6+2i)\]

OpenStudy (apoorvk):

Yeah, that is correct! Right-o @TramorrisM

OpenStudy (anonymous):

thanks ! (=